• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name A. Camsonne

    Log entry time 08:09:13 on June 2,2008

    Entry number 240763

    keyword=Owl shift summary

    June 2nd 2008
    SL : A. Camsonne
    TO : P. Decowski
    3rd : E. Piasetzky

    Kin 8 0% LH2 6 cm production




    0:00 taking over from swing shift run 1800

    00:03 Beam down because of RF problem

    00:11 Beam back

    1:05 ROC 10 crash

    1:39 Beam off for a few minutes

    6:15 Moving to C

    6:27

    6:32 Raster on

    6:46 spectrometer settled at 0.8948 GeV/c

    6:55 moved to LH2 for shower setup

    7:35 Guy is done setting up Bigbite shower starting production


    Run list production Kin 8 0%
    1800 5M
    1801 2.8 M
    1802 5M
    1803 5M
    1804 5M
    1805 5M
    1806 5M
    1807 5M

    68 M of events for this kinematic

    Spectrometer pointing runs
    1808 C raster off 264 k events
    1809 C raster on

    Bigbite shower setting run
    1810
    1811
    1812

    Bigbite production run
    1813 running